Christ is Risen (d minor)

Own Melody 12^{th} Century / Wittenberg 1529 / Choral Version: Norbert Südland 2010

[figure: 4_{2}; 4_{2}; 1. Christ; is; ri- -; sen; from; all; pain; and; suf-; fe-; ring;; there-; fore; shall; we; re-]

[figure: joice; and; sing,; Christ; brings; com-; fort; in; eve-; ry-; thing.; Ky-; ri-; e-; leis.]

[figure: 2. If; he; were; not; ri- -; sen,; the; world; would; be; in; pri- -; son;; since; he’s; ri-; sen]

[figure: from; the; pain,; we; praise; God; fa-; ther; not; in; vain.; Ky-; ri-; e-; leis.]

[figure: 3. Hal-; le- -; lu- -; jah,; hal-; le- -; lu- -; jah,; hal-; le- -; lu- -; jah!]

[figure: There-; fore; shall; we; re-; joice; and; sing,; Christ; brings; com-; fort; in; eve-; ry-; thing.; Ky-; ri-; e-; leis.]